OpenStudy (anonymous):

The angle of depression from the top of a flag pole to a point on the ground is 30°. If the point on the ground is 75 feet from the base of the flag pole, how tall is the pole to the nearest foot?

OpenStudy (anonymous):

does it mean this? 

OpenStudy (anonymous):

this

OpenStudy (anonymous):

ooooooooooh ok, then the ratios of your triangle are \(1:\sqrt{3}:2\) for short side : long side : hypotenuse

OpenStudy (anonymous):

since your long side is 75, the short side must be \(\frac{75}{\sqrt{3}}\)

OpenStudy (anonymous):

A) 38 feet B) 43 feet C) 53 feet D) 106 feet

OpenStudy (anonymous):

and so the height is \[\frac{75}{\sqrt{33}}\] or approximately 43.3

OpenStudy (anonymous):

ok i guess you should go with B, 43